# IntactKV: Improving Large Language Model Quantization by Keeping Pivot Tokens Intact

### v1

@@ 4.2 Main Results
We first integrate our proposed IntactKV with AWQ on LLaMA and LLaMA-2 models. The effect of this integration on model performance was measured by the perplexity (PPL) metric, with results on the C4 dataset detailed in Table 1, and results on the WikiText2 dataset in Appendix F.1. As indicated in these tables, IntactKV notably enhances the generative capabilities of AWQ-quantized models across various sizes, consistently surpassing the prior state-of-the-art (SOTA) method, OmniQuant. These findings demonstrate the efficacy of IntactKV in improving quantized LLMs and particularly highlight the effectiveness of utilizing the KV cache from full-precision models.
@@
#bits Method Vicuna-v1.5-7B Vicuna-v1.5-13B
FP16 - 5.31 5.52
w3g128 RTN 4.34 5.13
@@
  +IntactKV${}_{\textrm{[P]}}$+FT 4.84 5.44
w4g128 RTN 5.18 5.47
OmniQuant 5.09 5.48
AWQ 5.22 5.28
  +IntactKV${}_{\textrm{[P]}}$ 5.32 5.35
  +IntactKV${}_{\textrm{[P]}}$+FT 5.36 5.50
Table 5: GPT-4 evaluation of quantized Vicuna-v1.5 models on MT-Bench. The scores are on a scale of 10.
For SFT models, we implement IntactKV on the AWQ-quantized Vicuna models and evaluate the multi-task problem-solving ability on the MMLU benchmark. Table 2 presents the detailed zero-shot and five-shot results for Vicuna-v1.5-13B. The results demonstrate that IntactKV significantly enhances the performance of the AWQ-quantized model across all categories of tasks for Vicuna-v1.5-13B. Moreover, the performance of other model sizes under the five-shot setting is outlined in Table 3. Remarkably, AWQ+IntactKV exhibits superior performance over OmniQuant, achieving an average improvement of 1.09% across five model sizes. More results on MMLU are provided in Appendix F.2.
@@
#bits Method Vicuna-v1.5-7B Vicuna-v1.5-13B
FP16 - 5.31 5.52
w3g128 RTN 4.34 5.13
  +IntactKV${}_{\textrm{[P]}}$ 4.72 5.27
  +IntactKV${}_{\textrm{[P]}}$+FT 4.73 5.30
OmniQuant 4.78 5.05
  +IntactKV${}_{\textrm{[P]}}$ 4.94 5.10
  +IntactKV${}_{\textrm{[P]}}$+FT 4.85 5.24
AWQ 4.74 5.17
  +IntactKV${}_{\textrm{[P]}}$ 4.68 5.34
  +IntactKV${}_{\textrm{[P]}}$+FT 4.84 5.44
Table 6: Compatibility of IntactKV with various quantization methods. The scores are on a scale of 10.
Figure 4: Results of KV cache quantization with different bit-widths on 5-shot MMLU benchmark. Note that this is additional to INT3/4 weight quantization. Blue and red lines indicate quantizing model weights to INT3 and INT4, respectively. More experiment details can be found in Appendix G.
We further evaluate the quantized Vicuna models on zero-shot commonsense QA tasks. The results of Vicuna-v1.5-13B, as detailed in Table 4, indicate that IntactKV enables significant improvements over AWQ. Notably, our AWQ+IntactKV largely surpasses GPTQ and OmniQuant, especially under low-bit quantization. Additionally, Table 3 presents the average accuracy for various sizes of Vicuna models. In these evaluations, our AWQ+IntactKV consistently achieves the best zero-shot performance, which strongly demonstrates the efficacy of our proposed IntactKV. More results on Commonsense QA tasks can be found in Appendix F.3.
@@
To evaluate the quantized models’ generation capabilities in multi-turn conversations and their alignment with human preferences, we use GPT-4 to score the responses of quantized models on MT-Bench. We also fine-tune IntactKV, denoted as IntactKV+FT. From Table 5, IntactKV+FT significantly boosts the quantized model and consistently surpasses the GPTQ and OmniQuant for both INT3-g128 and INT4-g128 settings. For example, the 3-bit Vicuna-v1.5-13B quantized by AWQ has been improved from 5.17 to 5.34 by using the IntactKV, which can be further boosted to 5.44 with further fine-tuning. Remarkably, with trainable IntactKV, the setting of INT4-g128 even matches the full-precision model, while all other methods clearly lag behind the full-precision model. These results demonstrate the effectiveness of treating IntactKV as trainable parameters. Notably, the training process for the 7B model takes only 10 minutes on a single computing device, which is quite lightweight.
4.3 IntactKV with Other PTQ Methods
To assess IntactKV’s compatibility with different quantization methods, we build IntactKV upon various PTQ methods (i.e., RTN, OmniQuant, and AWQ) and evaluate on MT-bench with INT3-g128 quantization. From Table 6, IntactKV with different PTQ methods lead to an average boost of 0.16 in the final score for Vicuna-v1.5-7B, and 0.12 for Vicuna-v1.5-13B. Fine-tuning IntactKV further improves the performance of quantized models. Our IntactKV+FT consistently surpasses the performance of original quantized models for all three different quantization methods, raising the final score by 0.19 for Vicuna-v1.5-7B and 0.21 for Vicuna-v1.5-13B on average. These findings validate IntactKV’s capability of acting as a lightweight plugin, consistently improving various quantized models with negligible extra costs.
@@ 4.4 Extension to KV Cache Quantization
IntactKV can be readily applied to KV cache quantization to further decrease memory requirements. Similar to weight-only quantization, the quantized KV cache needs to be de-quantized before the matrix multiplication. Nonetheless, the proposed IntactKV is kept in FP16 since it is more sensitive to quantization. Also, the concatenation between the FP16 IntactKV and the rest de-quantized KV cache incurs negligible extra overhead. From Figure 4, IntactKV notably improves AWQ across different models and KV cache bit widths under the INT3 weight quantization. For INT4 weight quantization, IntactKV still gains an average accuracy increase of 0.27% over AWQ. We also notice that quantizing the KV cache to INT8 leads to almost no performance drop on the MMLU benchmark. When equipped with IntactKV, INT8 KV cache can even surpass vanilla AWQ-quantized models with FP16 KV cache, especially under INT3 weight quantization.
@@ 6 Limitations
Firstly, it is non-trivial to integrate IntactKV into activation quantization. For activation quantization, the whole KV cache needs to be quantized to low bits to exploit integer multiplications in self-attention, which contradicts our idea of keeping pivot tokens’ KV cache intact. Since IntactKV is treated as extra cached parameters, a straightforward solution is to quantize IntactKV with suitable quantization methods such as GPTQ. We leave this as future work. Secondly, more experiments may be needed for LLM evaluation. LLMs are being applied to a wide range of tasks, posing high demands on various model abilities. When quantizing LLMs to low bits, these abilities may be affected to varying degrees. Therefore, a comprehensive evaluation is required to gauge the capabilities of quantized LLMs. Although we experiment on several downstream tasks, such as PPL, MMLU, commonsense QA, and MT-bench, we note that this may not be enough to assess all abilities of LLMs. For example, how long context affects quantized models still remains unknown.
@@ Appendix A Proof of Theorem 1
FP16 5.69 5.08 4.09 3.52 5.12 4.57 3.12
RTN 6.98 5.88 4.84 4.22 6.21 5.17 3.74
GPTQ 6.62 5.68 4.75 4.20 6.68 5.14 3.79
OmniQuant 6.18 5.46 4.57 4.00 5.77 4.96 3.58
AWQ 6.34 5.53 4.60 3.95 5.82 4.97 3.53
  +IntactKV${}_{\textrm{[B]}}$ 6.23 5.49 4.54 3.89 5.73 4.95 3.50
Table 7: INT3-group128 quantization results of LLaMA and LLaMA-2 Models on WikiText2 dataset.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg. Hums STEM Social Others Avg.
OmniQuant 42.57% 35.98% 51.64% 53.55% 45.68% 42.95% 37.57% 53.59% 53.39% 46.54%
OmniQuant 49.52% 41.22% 59.47% 57.74% 51.82% 49.12% 44.40% 60.48% 58.95% 52.86%
OmniQuant 41.36% 33.96% 47.42% 47.69% 42.56% 41.15% 35.45% 48.20% 48.58% 43.18%
OmniQuant 43.29% 36.65% 51.64% 53.05% 45.95% 45.29% 37.84% 55.02% 54.38% 47.92%
OmniQuant 51.14% 42.08% 63.60% 59.84% 53.93% 53.77% 43.80% 63.47% 59.69% 55.12%
Table 8: INT3-group128 quantization results of Vicuna models on 0-shot and 5-shot MMLU benchmarks.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg. Hums STEM Social Others Avg.
OmniQuant 45.61% 38.70% 55.64% 56.23% 48.78% 45.21% 38.77% 57.17% 57.37% 49.25%
OmniQuant 50.01% 43.70% 62.33% 62.00% 54.12% 51.67% 43.87% 63.34% 61.81% 54.89%
OmniQuant 42.81% 34.72% 52.26% 52.00% 45.26% 43.21% 37.81% 52.62% 53.24% 46.43%
OmniQuant 47.93% 39.50% 57.98% 57.31% 50.48% 49.18% 39.86% 59.41% 58.14% 51.49%
OmniQuant 53.22% 44.73% 67.53% 63.05% 56.80% 56.68% 45.46% 68.67% 62.49% 58.24%
Table 9: INT4-group128 quantization results of Vicuna models on 0-shot and 5-shot MMLU benchmarks.
w3g128 RTN 39.40% 66.22% 43.34% 67.72% 77.40% 71.71% 61.17% 60.99%
GPTQ 37.80% 63.30% 41.81% 64.73% 74.46% 65.97% 58.57% 58.09%
OmniQuant 41.40% 66.22% 43.52% 67.26% 77.13% 70.35% 66.23% 61.73%
AWQ 40.80% 67.01% 43.34% 67.55% 78.32% 71.03% 64.97% 61.86%
w4g128 RTN 43.80% 68.82% 45.39% 70.33% 81.10% 73.31% 69.09% 64.55%
GPTQ 44.40% 69.93% 44.71% 70.45% 80.76% 73.72% 70.29% 64.89%
OmniQuant 43.00% 68.75% 44.28% 70.50% 81.01% 72.82% 70.10% 64.35%
AWQ 43.40% 68.75% 45.39% 70.66% 81.35% 73.43% 69.47% 64.64%
w3g128 RTN 42.00% 70.01% 47.44% 72.77% 82.20% 74.18% 69.20% 65.40%
GPTQ 42.40% 69.53% 48.46% 71.84% 83.76% 74.48% 70.00% 65.78%
OmniQuant 43.40% 68.51% 47.53% 71.09% 82.32% 73.92% 69.84% 65.23%
AWQ 44.00% 68.75% 48.21% 71.76% 83.58% 75.09% 70.68% 66.01%
w4g128 RTN 44.20% 70.80% 48.98% 73.82% 84.68% 76.36% 73.04% 67.41%
GPTQ 45.80% 70.96% 50.51% 73.99% 85.47% 76.70% 73.43% 68.12%
OmniQuant 43.80% 70.24% 49.74% 73.61% 84.59% 76.35% 72.54% 67.27%
AWQ 44.00% 72.06% 49.15% 73.44% 85.17% 77.00% 72.77% 67.66%
w3g128 RTN 41.00% 64.64% 39.08% 65.03% 75.81% 68.59% 59.73% 59.13%
GPTQ 39.60% 64.17% 42.24% 65.95% 72.63% 70.36% 63.65% 59.80%
OmniQuant 42.00% 67.88% 40.02% 66.75% 75.41% 70.52% 64.97% 61.08%
AWQ 41.60% 67.96% 38.65% 65.95% 76.36% 71.27% 64.76% 60.94%
w4g128 RTN 41.80% 67.80% 44.45% 69.91% 75.57% 73.25% 67.26% 62.86%
GPTQ 44.20% 68.75% 43.60% 70.88% 75.20% 73.39% 69.42% 63.63%
OmniQuant 42.40% 68.98% 44.03% 71.04% 76.97% 73.19% 69.49% 63.73%
AWQ 42.80% 67.09% 43.43% 71.34% 76.36% 73.46% 68.78% 63.32%
w3g128 RTN 42.20% 69.77% 44.03% 68.14% 80.52% 73.28% 64.68% 63.23%
GPTQ 42.00% 68.35% 44.88% 70.08% 80.67% 74.24% 68.00% 64.03%
OmniQuant 43.60% 70.56% 44.62% 70.66% 81.28% 74.13% 68.79% 64.81%
AWQ 42.60% 68.75% 45.90% 69.57% 80.80% 74.78% 69.30% 64.53%
w4g128 RTN 43.80% 71.19% 47.01% 72.64% 82.35% 76.19% 71.22% 66.34%
GPTQ 44.60% 70.01% 47.87% 73.32% 82.23% 76.55% 71.78% 66.62%
OmniQuant 45.80% 71.35% 46.08% 72.14% 82.35% 76.34% 71.38% 66.49%
AWQ 44.00% 70.01% 46.67% 72.64% 82.66% 76.43% 71.74% 66.31%
w3g128 RTN 46.00% 72.45% 49.83% 71.63% 82.75% 77.86% 69.82% 67.19%
GPTQ 44.40% 72.14% 47.61% 69.65% 83.49% 77.60% 71.86% 66.68%
OmniQuant 45.60% 73.32% 47.95% 72.52% 83.58% 78.05% 72.06% 67.58%
AWQ 45.40% 72.77% 50.77% 72.56% 82.42% 78.33% 71.45% 67.67%
w4g128 RTN 46.60% 73.95% 51.96% 74.07% 83.43% 80.01% 73.43% 69.06%
GPTQ 47.00% 73.48% 50.85% 73.06% 83.67% 80.31% 72.50% 68.70%
OmniQuant 47.60% 73.72% 51.96% 73.32% 83.58% 79.85% 73.57% 69.09%
AWQ 46.20% 73.40% 50.60% 73.53% 84.04% 79.91% 73.39% 68.72%
Table 10: Quantization results of Vicuna models on seven 0-shot commonsense QA tasks.
Model Download URL
LLaMA-2-7B https://huggingface.co/meta-llama/Llama-2-7b
LLaMA-2-13B https://huggingface.co/meta-llama/Llama-2-13b
LLaMA-2-70B https://huggingface.co/meta-llama/Llama-2-70b
Vicuna-v1.3-7B https://huggingface.co/lmsys/vicuna-7b-v1.3
Vicuna-v1.3-13B https://huggingface.co/lmsys/vicuna-13b-v1.3
Vicuna-v1.3-33B https://huggingface.co/lmsys/vicuna-33b-v1.3
Vicuna-v1.5-7B https://huggingface.co/lmsys/vicuna-7b-v1.5
Vicuna-v1.5-13B https://huggingface.co/lmsys/vicuna-13b-v1.5
Table 11: Download links to officially released LLMs.
@@ C.1 Implementation Details
We use ShareGPT dataset for our visualizations, where each sample starts with Vicuna system prompt of length 34. We use a randomly sampled sequence of length 128 to visualize the output activations and plot the corresponding attention map of the first 64 tokens. The attention score is mean pooled over different heads.
@@ Appendix D Experiment Details of Figure 2
We plot the quantization loss of the last Transformer layer as well as the total quantization loss of all attention layers with respect to the size of IntactKV on four different models, i.e., LLaMA-13B, LLaMA-30B, LLaMA-2-7B, and LLaMA-2-70B, covering different model types and model sizes. We use lossless IntactKV generated by the full-precision model to quantify the effect of IntactKV on the quantized model. IntactKV of size $s$ can ensure that the KV cache of the first $s$ tokens of the input sequence is generated by the full-precision model and thus lossless. Quantization loss is computed with MSE loss between the output activations of the quantized model and the full-precision model. We sample 128 sequences from the ShareGPT dataset to construct the validation set, each with a common prompt prefix of length 34. MSE loss is calculated on the tokens after the common prompt prefix. We quantize the model weights to 3 bits using round-to-nearest quantization with a group size of 128.
Appendix E Evaluation Details

### v2

@@ 4.2 Main Results
We first integrate our proposed IntactKV with RTN, GPTQ, OmniQuant, and AWQ on LLaMA and LLaMA-2 models. The effect of this integration on model accuracy is measured by the perplexity (PPL) metric, with results on the C4 dataset detailed in Table 1, and results on the WikiText2 dataset in Table 7. As indicated in these tables, IntactKV notably enhances the generative capabilities of quantized models across various LLMs and quantization methods, with AWQ+IntactKV  consistently achieving new state-of-the-art (SOTA) results. These findings demonstrate the efficacy of IntactKV in improving quantized LLMs and particularly highlight the effectiveness of utilizing the lossless KV cache from full-precision models. We provide more experiment results on LLaMA-3 and other heterogeneous LLMs (e.g. OPT) in Appendix G.1. IntactKV significantly improves different quantized LLMs, especially for LLaMA-3 models with larger quantization error. These results further prove the compatibility of our IntactKV with various LLM backbones.
@@
Method Vicuna-v1.5-7B Vicuna-v1.5-13B
FP16 5.31 5.52
RTN 4.34 5.13
    +IntactKV${}_{\textrm{[P]}}$ 4.72 5.27
    +IntactKV${}_{\textrm{[P]}}$+Cal 4.73 5.30
@@
    +IntactKV${}_{\textrm{[P]}}$+Cal 4.84 5.44
Table 5: GPT-4 evaluation of INT3-group128 weight-only quantized Vicuna-v1.5 models on MT-Bench. The scores are on a scale of 10.
Figure 4: Results of weight and KV cache quantization with different bit-widths on 5-shot MMLU benchmark. Note that this is additional to INT3/4 weight-only quantization. Blue and red lines indicate quantizing model weights to INT3 and INT4, respectively. We apply asymmetric per-head dynamic quantization to the KV cache.
Method LLaMA-7B LLaMA-13B LLaMA-2-7B LLaMA-2-13B LLaMA-3-8B
C4 WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2 C4 WikiText2
FP16 7.36 5.69 6.82 5.08 7.28 5.48 6.75 4.89 9.48 6.15
OmniQuant 17.03 12.17 15.65 11.16 21.40 14.74 16.24 12.28 - -
    +IntactKV${}_{\textrm{[B]}}$ 16.24 11.32 13.87 10.04 20.01 13.70 15.91 11.00 - -
QuaRot 8.23 6.29 7.40 5.55 8.30 6.11 7.51 5.39 13.42 8.21
    +IntactKV${}_{\textrm{[B]}}$ 8.05 6.15 7.32 5.45 8.12 5.97 7.25 5.21 12.23 7.54
Table 6: INT4 weight and activation quantization results of LLaMA models on C4 and WikiText2 datasets.
For SFT models, we implement IntactKV on the quantized Vicuna models and evaluate the multi-task problem-solving ability on the MMLU benchmark. Table 3 presents the detailed zero-shot and five-shot results for Vicuna-v1.3-13B. The results demonstrate that IntactKV significantly enhances the performance of quantized models across all categories of tasks and various quantization methods for Vicuna-v1.3-13B. Moreover, performance of Vicuna family under the five-shot setting is outlined in Table 2. Remarkably, IntactKV achieves an average improvement of 1.05% over OmniQuant and 0.8% over AWQ across five model sizes, with AWQ+IntactKV exhibiting superior performance over all the other quantized models. More results on MMLU are provided in Appendix G.2.
@@
We further evaluate the quantized Vicuna models on zero-shot commonsense QA tasks. The results of Vicuna-v1.3-13B, as detailed in Table 4, indicate that IntactKV enables significant improvements over various quantization methods. For example, AWQ+IntactKV surpasses the average accuracy of AWQ by 0.46% under INT3-g128 quantization. Additionally, Table 2 presents the average accuracy for various sizes of Vicuna models. In these evaluations, our IntactKV leads to an average accuracy improvement of 0.45% across different LLMs and quantization methods, which strongly demonstrates the efficacy of our proposed IntactKV. More results on commonsense QA tasks can be found in Appendix G.3.
@@
To evaluate the quantized models’ generation capabilities in multi-turn conversations and their alignment with human preferences, we use GPT-4 to score the responses of quantized models on MT-Bench. We also calibrate IntactKV, denoted as IntactKV+Cal. From Table 5, IntactKV significantly boosts the quantized model and IntactKV+Cal further enhances generation quality by compensating for the quantization error. For example, the 3-bit Vicuna-v1.5-13B quantized by AWQ has been improved from 5.17 to 5.34 by using the IntactKV, which can be further boosted to 5.44 with trainable IntactKV. We provide INT4 quantization results in Table 13. Remarkably, with trainable IntactKV, AWQ+IntactKV even matches the full-precision model under INT4 quantization, while all other methods clearly lag behind the full-precision model. These results demonstrate the effectiveness of IntactKV as well as treating IntactKV as trainable parameters. Notably, the training process for the 7B model takes only 10 minutes on a single computing device, which is quite lightweight. In Appendix H, we further demonstrate the effectiveness of calibrating IntactKV by comparing it with group bias tuning, a commonly used fine-tuning strategy for quantized models. IntactKV calibration can achieve better or comparable results with group bias tuning while using significantly fewer trainable parameters. Besides, IntactKV calibration serves as a more versatile calibration strategy for quantized models, which is suitable for various quantization settings.
@@ 4.3 Extension to KV Cache Quantization
IntactKV can be readily applied to KV cache quantization to further decrease memory requirements. We employ a mixed-precision strategy that keeps IntactKV in FP16 while the rest of the KV cache is quantized to lower bits. This only induces negligible memory overhead since IntactKV only contains the KV cache of the first few tokens. Note that this does not bring any additional inference costs since in the workflow of KV cache quantization, all quantized KV cache needs to be de-quantized back to FP16 before the matrix multiplication. Keeping IntactKV in FP16 reduces the overhead of de-quantization, i,e., we only need to cheaply concatenate the FP16 IntactKV with the rest de-quantized KV cache together. From Figure 4, IntactKV notably improves AWQ across different models and KV cache bit widths under the INT3 weight quantization. For INT4 weight quantization, AWQ+IntactKV still gains an average accuracy increase of 0.27% over the original quantized model. We also notice that quantizing the KV cache to INT8 leads to almost no performance drop on the MMLU benchmark. When equipped with IntactKV, INT8 KV cache can even surpass vanilla AWQ-quantized models with FP16 KV cache, especially under INT3 weight quantization.
4.4 Extension to Activation Quantization
In Table 6, we provide experiment results of combining IntactKV with OmniQuant Shao et al. (2024) and QuaRot Ashkboos et al. (2024) for weight and activation quantization. The implementation details can be found in Appendix E. To avoid extra inference costs, we need to quantize the whole KV cache to lower bits and can not keep the KV cache of pivot tokens intact. However, as detailed in Appendix I, we find that IntactKV has a significantly smoother distribution compared with the rest of the KV cache. Therefore, the full-precision IntactKV can be readily quantized to lower bits with negligible accuracy loss, thus rendering IntactKV amenable to weight and activation quantization with no extra inference costs. As shown in Table 6, our IntactKV significantly surpasses the performance of original quantized models for two different quantization methods, improving the PPL by 1.07 for OmniQuant and 0.31 for QuaRot on average. When combined with QuaRot, our IntactKV archives new state-of-the-art (SOTA) results on INT4 weight and activation quantization.
@@ 6 Limitations
More experiments may be needed for LLM evaluation. LLMs are being applied to a wide range of tasks, posing high demands on various model abilities. When quantizing LLMs to low bits, these abilities may be affected to varying degrees. Therefore, a comprehensive evaluation is required to gauge the capabilities of quantized LLMs. Although we experiment on several downstream tasks, such as PPL, MMLU, commonsense QA, and MT-bench, we note that this may not be enough to assess all abilities of LLMs. For example, how long context affects quantized models still remains unknown.
@@ C.1 Implementation Details
We use ShareGPT dataset for our visualizations, where each sample starts with Vicuna system prompt of length 34. We use a randomly sampled sequence of length 128 to visualize the output activations and plot the corresponding attention map of the first 64 tokens for better visualization. The attention score is mean pooled over different heads.
@@ C.3 Visualization of Vicuna Models
C.4 Visualization of OPT and Mistral Models
To demonstrate the prevalence of pivot tokens in LLMs, we provide more visualizations on OPT and Mistral models in Figure. 2021. The results show that the pivot tokens with extreme outliers are ubiquitous in various LLMs.
@@ Appendix D Experiment Details of Figure 2
We plot the quantization loss of the last Transformer layer as well as the total quantization loss of all attention layers with respect to the size of IntactKV on four different models, i.e., LLaMA-13B, LLaMA-30B, LLaMA-2-7B, and LLaMA-2-70B, covering different model types and model sizes. We use lossless IntactKV generated by the full-precision model to quantify the effect of IntactKV on the quantized model. IntactKV of size $s$ can ensure that the KV cache of the first $s$ tokens of the input sequence are generated by the full-precision model and thus lossless. Quantization loss is computed with MSE loss between the output activations of the quantized model and the full-precision model. We sample 128 sequences from the ShareGPT dataset to construct the validation set, each with a common prompt prefix of length 34. MSE loss is calculated on the tokens after the common prompt prefix. We quantize the model weights to 3 bits using round-to-nearest quantization with a group size of 128.
Appendix E Quantization Method Details
We carefully reproduce the results of various quantization methods with their official code or released checkpoint.
Weight-only Quantization.
For GPTQ, we use AutoGPTQ [^2] https://github.com/AutoGPTQ/AutoGPTQ with C4 calibration set following (Frantar et al., 2022) to reproduce all results. We turn the quantization option "–desc_act" on to quantize weight columns in order of decreasing activation size, which is a heuristic rule empirically found to be effective for GPTQ. For AWQ Lin et al. (2023), we directly load the officially released quantization parameters of LLaMA models for evaluation and reproduce results on Vicuna models with their official code [^3] https://github.com/mit-han-lab/llm-awq using Pile (Gao et al., 2020) calibration set. For weight-only quantization of OmniQuant, we reproduce results with their official code [^4] https://github.com/OpenGVLab/OmniQuant using WikiText2 (Merity et al., 2016) calibration set. We only activate the option "–lwc" to learn the weight clipping parameters for both LLaMA and Vicuna models, following (Shao et al., 2024). Additionally, for OmniQuant+IntactKV${}_{\textrm{[B]}}$, we directly integrate IntactKV${}_{\textrm{[B]}}$ into the training process of OmniQuant to adapt the weight clipping parameters to IntactKV${}_{\textrm{[B]}}$, which is found to be effective and introduces no extra training costs.
Method OPT-6.7B Mistral-7B
C4 WikiText2 C4 WikiText2
FP16 12.75 10.83 8.39 5.30
RTN 36.18 23.91 9.65 6.20
AWQ 13.39 11.38 9.29 5.95
    +IntactKV${}_{\textrm{[B]}}$ 13.37 11.32 9.25 5.93
Table 9: INT3-group128 weight-only quantization results of OPT and Mistral on C4 and WikiText2 datasets.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg Hums STEM Social Others Avg
OmniQuant 42.72% 36.38% 51.93% 53.55% 45.88% 42.70% 37.97% 54.31% 53.08% 46.62%
OmniQuant 49.73% 41.02% 59.31% 58.33% 51.94% 49.18% 44.17% 60.45% 58.91% 52.82%
OmniQuant 41.40% 34.06% 48.07% 48.06% 42.82% 40.98% 35.19% 48.23% 48.03% 42.95%
OmniQuant 43.51% 36.85% 52.16% 53.05% 46.18% 45.91% 37.44% 55.31% 54.94% 48.23%
OmniQuant 51.22% 42.18% 64.06% 60.39% 54.21% 53.94% 44.10% 63.21% 59.81% 55.21%
Table 10: INT3-group128 weight-only quantization results of Vicuna models on MMLU benchmarks.
Model Method MMLU (0 shot) MMLU (5 shot)
Hums STEM Social Others Avg Hums STEM Social Others Avg
OmniQuant 44.97% 38.80% 55.57% 56.32% 48.59% 45.53% 39.40% 57.20% 57.50% 49.53%
OmniQuant 49.99% 43.97% 62.40% 62.03% 54.19% 51.67% 43.90% 63.05% 61.81% 54.84%
OmniQuant 43.12% 34.59% 52.45% 52.31% 45.46% 43.04% 37.67% 52.75% 53.08% 46.33%
OmniQuant 47.52% 39.40% 57.98% 57.37% 50.34% 49.03% 40.09% 59.34% 58.11% 51.47%
OmniQuant 53.22% 44.43% 67.73% 63.26% 56.73% 56.83% 45.46% 68.67% 62.31% 58.25%
Table 11: INT4-group128 weight-only quantization results of Vicuna models on MMLU benchmarks.
w3g128 RTN 40.60% 66.22% 43.77% 67.89% 77.86% 71.46% 61.75% 61.36%
GPTQ 39.40% 64.72% 40.87% 65.07% 74.77% 66.32% 59.09% 58.61%
OmniQuant 43.00% 66.46% 43.69% 67.72% 78.59% 70.53% 66.12% 62.30%
AWQ 41.60% 67.56% 42.66% 67.85% 78.96% 71.32% 65.28% 62.18%
w4g128 RTN 43.40% 68.98% 44.80% 71.09% 82.05% 73.32% 69.28% 64.70%
GPTQ 43.60% 69.77% 44.62% 70.20% 74.01% 72.61% 68.27% 63.30%
OmniQuant 43.40% 69.06% 44.37% 71.17% 81.83% 72.90% 70.13% 64.69%
AWQ 43.80% 68.59% 45.73% 71.09% 82.02% 73.51% 69.42% 64.88%
w3g128 RTN 43.60% 71.27% 48.55% 72.81% 82.91% 74.55% 69.18% 66.12%
GPTQ 43.00% 70.09% 48.98% 72.98% 84.43% 74.80% 70.11% 66.34%
OmniQuant 43.60% 69.85% 47.78% 71.17% 82.45% 74.16% 70.04% 65.58%
AWQ 45.40% 69.38% 48.38% 71.89% 84.46% 75.24% 70.85% 66.51%
w4g128 RTN 44.80% 71.51% 49.15% 73.78% 85.20% 76.70% 72.62% 67.68%
GPTQ 45.80% 70.96% 50.51% 73.99% 85.47% 76.70% 73.43% 68.12%
OmniQuant 44.40% 70.80% 50.09% 73.86% 85.29% 76.79% 72.39% 67.66%
AWQ 45.60% 72.85% 49.49% 74.07% 85.72% 77.37% 72.37% 68.21%
w3g128 RTN 41.80% 63.38% 38.91% 63.47% 76.57% 68.92% 60.29% 59.05%
GPTQ 40.00% 65.90% 41.55% 66.16% 70.73% 69.66% 62.95% 59.56%
OmniQuant 42.00% 66.06% 39.68% 66.67% 75.69% 70.45% 65.65% 60.89%
AWQ 42.40% 66.69% 39.51% 65.40% 77.06% 70.53% 63.69% 60.75%
w4g128 RTN 42.20% 67.80% 43.00% 70.66% 75.50% 73.16% 68.37% 62.96%
GPTQ 45.20% 68.82% 42.41% 70.45% 67.58% 72.50% 67.40% 62.05%
OmniQuant 43.40% 67.96% 44.28% 71.46% 76.42% 73.22% 68.81% 63.65%
AWQ 43.60% 68.03% 43.26% 71.68% 75.87% 73.44% 68.45% 63.48%
w3g128 RTN 44.00% 70.96% 44.03% 67.30% 80.40% 73.33% 64.00% 63.43%
GPTQ 45.20% 69.77% 46.08% 70.33% 81.90% 74.89% 67.59% 65.11%
OmniQuant 45.20% 69.22% 45.22% 68.90% 80.95% 74.72% 68.15% 64.62%
AWQ 42.80% 68.98% 46.08% 68.98% 81.31% 74.97% 68.78% 64.56%
w4g128 RTN 45.20% 71.43% 48.04% 73.15% 82.87% 76.56% 70.62% 66.84%
GPTQ 44.60% 70.01% 47.87% 73.32% 82.23% 76.55% 71.78% 66.62%
OmniQuant 45.60% 70.56% 46.76% 73.02% 82.81% 76.74% 70.41% 66.56%
AWQ 45.20% 70.32% 47.27% 73.91% 82.81% 76.79% 71.32% 66.80%
w3g128 RTN 46.60% 72.53% 49.06% 72.18% 83.12% 78.06% 69.73% 67.33%
GPTQ 44.80% 71.74% 47.01% 70.12% 83.64% 77.79% 71.51% 66.66%
OmniQuant 45.40% 73.64% 48.63% 72.35% 83.55% 77.98% 71.73% 67.61%
AWQ 45.60% 73.32% 50.68% 71.63% 82.39% 78.55% 71.49% 67.67%
w4g128 RTN 47.20% 73.88% 51.62% 74.12% 83.58% 79.86% 73.24% 69.07%
GPTQ 47.00% 73.48% 50.85% 73.06% 83.67% 80.31% 72.50% 68.70%
OmniQuant 48.80% 74.19% 50.68% 73.91% 83.79% 79.83% 73.28% 69.21%
AWQ 47.00% 73.16% 50.85% 73.82% 84.19% 79.77% 73.32% 68.87%
Table 12: Weight-only quantization results of Vicuna models on seven 0-shot commonsense QA tasks.
Method Vicuna-v1.5-7B Vicuna-v1.5-13B
FP16 5.31 5.52
RTN 5.18 5.47
OmniQuant 5.09 5.48
AWQ 5.22 5.28
    +IntactKV${}_{\textrm{[P]}}$ 5.32 5.35
    +IntactKV${}_{\textrm{[P]}}$+Cal 5.36 5.50
Table 13: GPT-4 evaluation of INT4-group128 weight-only quantized Vicuna-v1.5 models on MT-Bench. The scores are on a scale of 10.
Weight and Activation Quantization.
For weight and activation quantization of OmniQuant, it is difficult to integrate IntactKV${}_{\textrm{[B]}}$ into training with the learnable equivalent transformation, so we reuse the official checkpoint of LLaMA and LLaMA-2 models. When combining IntactKV with OmniQuant, we quantize IntactKV to lower bits to avoid additional inference overhead. We do not include OmniQuant results on LLaMA-3 models since the option "–let" is not compatible with GQA (Group Query Attention). For QuaRot Ashkboos et al. (2024), we reproduce all the results with their official code [^5] https://github.com/spcl/QuaRot using WikiText2 (Merity et al., 2016) calibration set. We do not quantize IntactKV to lower bits since QuaRot adopts a mixed-precision self-attention quantization strategy and can not utilize the integer multiplications for self-attention operations. Therefore, maintaining IntactKV in FP16 will not bring any extra inference costs for QuaRot.
Appendix F Evaluation Details